OpenStudy (anonymous):

Square root of (4x^6)

OpenStudy (amistre64):

sqrt(4) * x^(6/2)

OpenStudy (jamesj):

No, not zero.

OpenStudy (jamesj):

Simplify what amistre has written

OpenStudy (anonymous):

Why did he divide the 6 by 2..?

OpenStudy (amistre64):

those root radicals are just exponents

OpenStudy (amistre64):

2rt() = ^1/2 10rt() = ^1/10

OpenStudy (jamesj):

because \[ (x^a)^b = x^{ab} \] Hence \[ \sqrt{x^6} = (x^6)^{1/2} = x^{6/2} = ... \]

OpenStudy (amistre64):

its makes the math simpler

OpenStudy (jamesj):

Hence here \[ \sqrt{4x^6} = \sqrt{4} \sqrt{x^6} = ... what? \]

OpenStudy (anonymous):

I still don't get where the 1/2 is coming from

OpenStudy (jamesj):

\[ \sqrt{x} = x^{1/2} \] For example, if \[ \sqrt{x} = 3 \] then \[ \sqrt{x}^2 = 3^2 \] i.e., \[ x = 3^2 \] It must be therefore that \( \sqrt{x} = x^{1/2} \). If this weren't the case we wouldn't have \[ \sqrt{x}^2 = x \] Having \[ \sqrt{x} = x^{1/2} \] is consistent with the definition of square root because \[ (x^{1/2})^2 = x^{2/2} = x^1 = x \]

OpenStudy (anonymous):

So am i only dividing the x^6 by two to make it x^3 since the exponent doesn't pertain to the 4?